The shaft of the tibia is subjected to compressive forces. The length of the tibia shaft is 0.30 m and the cross-sectional area of cortical bone in the shaft is 0.0002 m2. If the ultimate strength of cortical bone is 200 x 10^6 N/m2 and if the elastic modulus of the shaft of the tibia is 2 x 10^10 N/m2, what is the maximum compressive force (in N) that can be applied to the tibia shaft before compressive failure?

Now assume that the maximum force value you found above is applied as a tensile force instead of a compressive force. The tibia shaft would be _________ to experience failure.

A. Less likely

B. Equally likely

C. More likely
Answer #1

1)
ultimate strength = Fmax/A

==> Fmax = ultimate strength*A

= 200*10^6*0.0002

= 4.0*10^4 N

2) B. Equally likely

because, for a given material the effect due to tensile stress and comressive stress are same.